# Capturing Knowledge Graphs and Rules with Octagon Embeddings

### v1

@@ B.3 Capturing Non-Composition Rules
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ B.4 Capturing Rules: Compositions
[…]
@@ C.1 Implementation Details
The octagon model was implemented using the PyKEEN library [^5] https://github.com/pykeen/pykeen. Octagon embeddings are learnt in batches using the Adam optimizer and a Self-Adversarial Negative Sampling (NSSA) loss, like BoxE and ExpressivE. Validation is performed every 10 epochs and the best model over 1,000 epochs is used for testing. In all experiments, octagons are initialised as narrow $u$-bands with wide $v$, $x$ and $y$ constraints. That is, $(u_{w})_{i}=0.1$ and $(v_{w})_{i}=(x_{w})_{i}=(y_{w})_{i}=1$ for all $1\leq i\leq n$. The center of $u$-bands is initialised randomly.
@@
The full implementation, including configuration files to reproduce results discussed in Section C.2, is available online [^6] https://github.com/vcharpenay/uvxy. Hyperparameters used in the experiments are also reproduced in Table 2 for completeness. Some hyperparameters (embedding dimension, number of negative samples and batch size) alter the inductive power of a model. For a fair comparison, in high dimension, we have set these hyperparameters with the same values as in prior experiments on ExpressivE [^7] https://github.com/AleksVap/ExpressivE/, with one exception: on WN18RR, we empirically observed that less negative samples gave slightly better results. We therefore set $|N|=5$ (instead of $|N|=100$). In low dimension, these hyperparameters have been taken from experiments done on MuRE [^8] https://github.com/ibalazevic/multirelational-poincare/, with the same exception on WN18RR. To our knowledge, no result had been made available for BoxE and ExpressivE in low dimension. Other hyperparameters that are rather influenced by the model’s scoring function (margin, $L_{p}$-norm, learning rate) have been chosen after a grid search. Possible values were $\{3,4,6,9,12,15,21,27,35,41\}$ for the margin, $\{1,2\}$ for the $L_{p}$-norm and $\{10^{-2},10^{-3}\}$ for the optimizer’s learning rate. In low dimension, a search over $\{256,512,1024\}$ has also been performed for the optimal batch size.
@@ C.2 Discussion
$u$ 22.8 37.2 53.2 33.0 00.4 40.4 52.0 21.6
$ux$ 22.9 36.4 52.6 32.7 00.4 40.4 53.2 22.0
$uxy$ 22.8 36.5 52.3 32.6 01.7 46.7 53.8 22.5
$uv$ 23.0 35.7 51.1 32.3 01.9 42.4 52.1 23.2
$uv^{*}$ 23.1 36.3 52.3 32.7 42.8 46.5 50.1 45.5
$uvxy$ 22.8 35.6 50.6 32.0 04.8 43.5 53.4 25.5
$uvxy^{*}$ 23.1 36.0 52.1 32.6 44.3 49.9 56.6 48.5
$u+v$ 23.6 37.4 53.3 33.5 41.3 45.8 50.2 44.5
Table 3: Link prediction performances of region based embedding models ($n=1000$ on FB15k-237, $n=500$ on WN18RR). Configurations with $*{}^{*}$ use the variant with attention weights.
Table 3 gives results in high dimension. Results for TransE, BoxE and ExpressivE are not ours. They were respectively published in Sun et al. (2019)Abboud et al. (2020) and Pavlovic and Sallinger (2023). Two results are available for the $uv$ and $uvxy$ configurations: with and without attention weights. Attention improves performance on all metrics but one: Hits@10 on WN18RR in the $uv$ configuration is $2\%$ lower with attention. Attention does however greatly increase Hits@1.
@@
$u$ 19.9 32.7 48.6 29.4 02.0 34.4 49.7 20.3
$ux$ 20.3 32.9 48.6 29.7 02.1 34.8 50.2 20.6
$uy$ 20.1 32.9 48.6 29.5 02.0 32.8 49.4 19.7
$uxy$ 20.5 33.3 48.9 29.9 02.3 34.2 50.1 20.5
$uv$ 15.9 27.4 41.9 24.6 16.4 43.3 52.8 31.2
$uv^{*}$ 19.0 30.7 45.6 27.8 16.4 43.3 52.8 31.2
$uvxy$ 20.2 32.5 47.8 29.4 32.1 45.5 52.1 40.1
$uvxy^{\text{\textdagger}}$ 20.2 32.0 47.6 29.2 37.3 43.9 49.3 41.7
$uvxy^{*}$ 20.6 32.5 47.2 29.5 31.7 46.0 52.2 40.1
$u+v$ 19.9 32.0 47.9 29.0 30.7 41.9 49.9 37.6
Table 4: Link prediction of region based embedding models in low dimension ($n=40$). Configurations with $*{}^{*}$ use the variant with attention weights. Configurations with ${}^{\text{\textdagger}}$ use regular octagons at initialisation (others use pseudo $u$-bands).
We evaluated more configurations in low dimension. Results for TransE, BoxE and ExpressivE are also ours, those for MuRE were first published in Balazevic et al. (2019a). Table 4 gives results for 40-dimensional models. Several observations can be made. For instance, the performance of $uv$ and $uvxy$ in low dimension is comparable to their performance in higher dimension. We also observe that attention weights do not alter the performance of a model in low dimension, contrary to higher-dimensional models. Conversely, full $uvxy$-octagons outperform TransE, BoxE and ExpressivE on FB15k-237 in low dimension but not in high dimension. From Table 4, we also note that $ux$ and $uy$ give comparable results, with a small penalty for $uy$.
Two further results open up perspectives on learning octagon models. On WN18RR, a different initialisation of octagons has a strong impact on the learned embeddings: if initial octagons are regular octagons (instead of pseudo $u$-bands), Hits@1 increases by $5\%$, at the expense of lower Hits@3 and Hits@10. More elaborate initialisation techniques may better take advantage of the expressivity of octagons. Still on WN18RR, a model that mixes a $u$ constraint in half of the coordinates and a $v$ constraint in the other half achieves much higher Hits@1 than both $u$ and $uv$. This model is denoted $u+v$ in Table 4. If we compare $u+v$ with BoxE, we get opposite conclusions in low and high dimension. In low dimension, it has lower Hits@1 than BoxE but comparable Hits@10; in high dimension, it has higher Hits@1 but lower Hits@10. A more complex mixture of constraints may yield better performances.

### v2

@@ B.3 Capturing Non-Composition Rules
Note that $X_{j}^{\textit{inv}}=Y_{j}$. Let $O_{r,i}$ be the octagon representing relation $r$ in coordinate $i$. We initialise these octagons as follows:
@@
We now consider the inductive case. Suppose $X_{2}\subseteq O^{(j)}_{t,i}$. Note that all octagons are closures of subsets of $X_{1},X_{2},Y_{1},Y_{2}$. For such octagons, we have $X_{2}\subseteq\textit{cl}(O_{1},\ldots,O_{k})$ iff $X_{2}\subseteq O_{j}$ for some $j$. We thus have that one of the following cases must be true.
@@
Assume $\mathcal{K}\not\models r\subseteq s^{-1}$. There exists a coordinate $i$ such that $O_{r,i}\not\subseteq O_{s,i}^{\textit{inv}}$.
@@
We then again have that for any relation $t$ it holds that $X_{2}\subseteq O^{(j)}_{t,i}$ implies $\mathcal{K}\models r\subseteq t$ and $Y_{2}\subseteq O^{(j)}_{t,i}$ implies $\mathcal{K}\models r^{-1}\subseteq t$. We have in particular that $X_{2}\subseteq O^{(j)}_{s^{-1},i}$ iff $Y_{2}\subseteq O^{(j)}_{s,i}$. The latter implies $\mathcal{K}\models r^{-1}\subseteq s$, which is equivalent to $\mathcal{K}\models r\subseteq s^{-1}$. Since we assumed $\mathcal{K}\models r\not\subseteq s^{-1}$, it follows that $X_{2}\not\subseteq O^{(j)}_{s^{-1},i}$ and thus $O_{r,i}\not\subseteq O_{s,i}^{\textit{inv}}$. ∎
@@
Let $r\in\mathcal{R}$. For every coordinate $i$, it holds that $O_{r,i}\cap O_{r,i}^{\textit{inv}}\neq\emptyset$.
@@ B.4 Capturing Rules: Compositions
where the closure cl of a set of octagons $O_{1},\ldots,O_{k}$ with $O_{i}=\mathsf{Octa}(x_{i}^{-},\allowbreak x_{i}^{+},\allowbreak y_{i}^{-},\allowbreak y_{i}^{+},\allowbreak u_{i}^{-},\allowbreak u_{i}^{+},\allowbreak v_{i}^{-},\allowbreak v_{i}^{+})$ is given by the octagon $O^{*}=\mathsf{Octa}(x_{*}^{-},\allowbreak x_{*}^{+},\allowbreak y_{*}^{-},\allowbreak y_{*}^{+},\allowbreak u_{*}^{-},\allowbreak u_{*}^{+},\allowbreak v_{*}^{-},\allowbreak v_{*}^{+})$ with $x_{*}^{-}=\min_{i}x_{i}^{-}$, $x_{*}^{+}=\max_{i}x_{i}^{+}$, $y_{*}^{-}=\min_{i}y_{i}^{-}$, $y_{*}^{+}=\max_{i}y_{i}^{+}$, $u_{*}^{-}=\min_{i}u_{i}^{-}$, $u_{*}^{+}=\max_{i}u_{i}^{+}$, $v_{*}^{-}=\min_{i}v_{i}^{-}$ and $v_{*}^{+}=\max_{i}v_{i}^{+}$. In other words, $\textit{cl}(O_{1},\ldots,O_{k})$ is the smallest octagon which contains $O_{1},\ldots,O_{k}$.
@@ C.1 Implementation Details
The octagon model was implemented using the PyKEEN library [^7] https://github.com/pykeen/pykeen. Octagon embeddings are learnt in batches using the Adam optimizer and a Self-Adversarial Negative Sampling (NSSA) loss, like BoxE and ExpressivE. Validation is performed every 10 epochs and the best model over 1,000 epochs is used for testing. In all experiments, octagons are initialised randomly with Xavier initialisation. We empirically observed that normalisation (Proposition 11) does not influence final performances. Octagons are thus unconstrained during training, for faster computation.
@@
The full implementation, including configuration files to reproduce results discussed in Section C.2, is available online [^8] https://github.com/vcharpenay/uvxy. Hyperparameters used in the experiments are also reproduced in Table 2 for completeness. Some hyperparameters (embedding dimension, number of negative samples and batch size) alter the inductive power of a model. For a fair comparison, we chose the same configuration as past experiments for these hyperparameters. In high dimension, we took ExpressivE experiments [^9] https://github.com/AleksVap/ExpressivE/ as a basis, up to one exception: on WN18RR, we empirically observed that less negative samples gave slightly better results. We therefore set $|N|=5$ instead of $|N|=100$. In low dimension, we took hyperparameters used on MuRE [^10] https://github.com/ibalazevic/multirelational-poincare/, with the same exception on WN18RR. To our knowledge, no result had been made available for BoxE and ExpressivE in low dimension. Other hyperparameters that are rather influenced by the model’s scoring function (margin, $L_{p}$-norm, learning rate) have been chosen after a grid search. Possible values were $\{3,4,6,9,12,15,21,27,35,41\}$ for the margin, $\{1,2\}$ for the $L_{p}$-norm and $\{10^{-2},10^{-3}\}$ for the optimizer’s learning rate. In low dimension, a search over $\{256,512,1024\}$ has also been performed for the optimal batch size.
@@ C.2 Discussion
$u$ 23.1 37.3 53.2 33.2 01.6 39.9 51.5 22.0
$ux$ 23.3 37.1 52.5 33.1 01.9 39.0 51.6 21.8
$uxy$ 23.2 37.2 53.1 33.2 01.8 40.8 52.8 22.8
$uv$ 22.7 35.8 51.2 32.2 40.4 46.2 51.3 44.4
$uv^{*}$ 24.1 36.9 52.8 33.5 43.6 48.5 52.9 46.9
$uvxy$ 15.9 28.0 42.9 24.9 39.6 45.9 50.7 43.8
$uvxy^{*}$ 24.1 36.7 51.7 33.2 43.6 49.2 56.1 47.9
$u+v$ 24.0 37.2 53.0 33.6 42.4 46.2 51.8 45.5
Table 3: Link prediction performances of region based embedding models ($n=1000$ on FB15k-237, $n=500$ on WN18RR). Configurations with use the variant with attention weights.
Table 3 gives results in high dimension. Results for TransE, BoxE and ExpressivE are not ours. They were respectively published in Sun et al. (2019)Abboud et al. (2020) and Pavlovic and Sallinger (2023). Two results are available for the $uv$ and $uvxy$ configurations: with and without attention weights. Attention improves performance on all metrics. In the light of Propositions 5 and  16, which highlight the importance of $v$ constraints, we also introduced a $u+v$ model consisting of a mixture a $u$ constraint in half of the coordinates and a $v$ constraint in the other half. This model clearly outperforms $u$ on WN18RR, especially on H@1, though remaining below $uv^{*}$ and $uvxy^{*}$ (with attention weights).
@@
$u$ 20.3 32.8 48.9 29.6 02.4 36.9 49.1 21.0
$u^{*}$ 18.1 29.3 43.3 26.6 22.2 40.1 51.5 32.9
$ux$ 20.8 32.9 48.5 30.0 02.6 36.3 49.8 21.0
$uy$ 20.3 32.8 48.5 29.6 02.5 38.0 49.8 21.6
$uxy$ 20.5 32.9 48.5 29.7 02.4 37.1 49.0 21.1
$uv$ 20.3 32.3 47.5 29.3 38.5 46.5 52.3 43.6
$uv^{*}$ 19.6 31.1 45.1 28.2 41.1 47.0 51.1 44.9
$uvxy$ 19.7 31.5 46.8 28.6 37.9 44.6 48.9 42.1
$uvxy^{*}$ 20.4 31.6 46.3 28.9 35.6 43.8 49.1 40.6
$u+v$ 20.2 32.3 48.0 29.3 40.9 46.3 50.5 44.5
Table 4: Link prediction of region based embedding models in low dimension ($n=40$). Configurations with use the variant with attention weights.
We evaluated more configurations in low dimension, as reported in Table 4. Results for MuRE were first published in Balazevic et al. (2019a), others are ours. From the table, we observe that attention weights do not always improve the performances of a model in low dimension contrary to higher-dimensional models. Conversely, most models derived from $uvxy$ outperform TransE, BoxE and ExpressivE on FB15k-237 in low dimension but not in high dimension. We also note that $ux$ and $uy$ give comparable results, with a small penalty for $uy$.
Acknowledgments